OpenStudy (anonymous):

Two cardboard boxes full of books are in contact with each other on a friction less table. Box H has twice the mass of box G. If you push on a box G with a horizontal force F, then box H will experience a net force of what?

OpenStudy (shamim):

i think the box h will experience a force equal to f

OpenStudy (shamim):

if u need explanation then i wanna tell the whole story

OpenStudy (shamim):

i think u r now offline

OpenStudy (anonymous):

For these type of questions, you'll need Newton's second and third law. as they are connected, the acceleration they experience is the same. Let mass of G be m and that of H be 2m. total acceleration: \(F=3ma\) \(a=\frac{F}{3m}\)  For H, [f is the net force] \(f=2ma\) \(f=2m(\frac{F}{3m})\) \(f=\frac{2F}{3}\) As such the net force on H is two-thirds of F.